# T-JEPA: Augmentation-Free Self-Supervised Learning for Tabular Data

### v1

@@ Appendix B Compute and training time
AD HI HE JA AL CA
Samples 48,842 98,050 65,196 83,733 108,000 20,640
Numerical, Categorical 66, 88 $28$, 0 $27$, 0 $54$, 0 $128$, 0 88, 0
Classes 22 22 $100$ 44 $1,000$ N/A
Metric Accuracy Accuracy Accuracy Accuracy Accuracy RMSE
Pretraining GPU-hours $0.84$ $1.80$ $1.03$ $1.27$ $3.64$ $0.34$
Appendix C Alternative Strategies
Other settings have been considered before the one discussed in section 3, and were discarded because they led to collapsed regimes. We considered two alternatives. First, we considered a pipeline where only the masking strategy differs, as detailed in section C.1. Second, we also considered an alternative where the masking strategy is the one detailed in section C.1, and we also modify the predictor architecture as detailed in section C.2.
C.1 Masking Strategy
@@
Given this modification, the obtained context representation’s dimension differ from the one given in equation (2).
$h_{\text{context}}^{\mathbf{m}}=f_{\theta}(\mathbf{z}_{\mathbf{x}}^{\mathbf{m}})\in\mathbb{R}^{d\times h}\quad(\text{context})$
The rest of the pipeline is identical to the one given in section 3. We also considered a second alternative where the above pipeline is chosen but a different predictor architecture (see section C.2) replaces the transformer as detailed in section 3.
C.2 Predictor
@@
Appendix D Embedding feature variance
@@
Appendix E Detailed Performance
Table 6: Performance metrics for different models and T-JEPA enhancement across various datasets including the standard deviations. The metric for models without T-JEPA and augmented by PTaRL are taken from [Ye et al., 2024] in which no standard deviations were reported. We also include for comparison SSL methods like SwitchTab [Wu et al., 2024], BinRecon [Lee et al., 2024] or VIME [Yoon et al., 2020]. We report in bold the metric that wins between the raw data representation and the augmented representations. We report in blue the best performing deep method (including SSL methods) and underline the overall best metric for a dataset.
AD $\uparrow$ HE $\uparrow$ JA $\uparrow$ AL $\uparrow$ CA $\downarrow$ HI $\uparrow$ Wins Baseline Neural Networks MLP $0.825$ $0.352$ $0.672$ $0.917$ $0.518$ $0.681$ 0 +PTaRL $\mathbf{0.868}$ $0.396$ $0.710$ $\mathbf{0.964}$ $0.489$ $\mathbf{0.723}$ 33 +T-JEPA $0.866$ $\pm 2e^{-3}$ $\mathbf{0.400}$ $\pm 4e^{-3}$ $\mathbf{0.728}$ $\pm 3e^{-3}$ $0.961$ $\pm 6e^{-3}$ $\mathbf{0.468}$ $\pm 3.7e^{-2}$ $0.517$ $\pm 8e^{-2}$ 33 DCNv2 $0.826$ $0.340$ $0.662$ $0.905$ $0.502$ $0.681$ 0 +PTaRL $\mathbf{0.867}$ $0.389$ $\mathbf{0.723}$ $\mathbf{0.959}$ $0.465$ $\mathbf{0.731}$ 33 +T-JEPA $0.861$ $\pm 2e^{-3}$ $\mathbf{0.399}$ $\pm 3e^{-3}$ $\mathbf{0.723}$ $\pm 2e^{-3}$ $0.955$ $\pm 2e^{-3}$ $\mathbf{0.420}$ $\pm 2.6e^{-2}$ $0.525$ $\pm 8e^{-2}$ 33 ResNet $0.813$ $0.354$ $0.666$ $0.919$ $0.537$ $0.682$ 0 +PTaRL $0.862$ $0.399$ $\mathbf{0.723}$ $\mathbf{0.964}$ $0.498$ $\mathbf{0.729}$ 22 +T-JEPA $\mathbf{0.865}$ $\pm 3e^{-3}$ $\mathbf{0.401}$ $\pm 2e^{-3}$ $0.718$ $\pm 3e^{-3}$ $\mathbf{0.964}$ $\pm 1e^{-3}$ $\mathbf{0.441}$ $\pm 8e^{-2}$ $0.705$ $\pm 5e^{-3}$ 44 AutoInt $0.823$ $0.338$ $0.653$ $0.894$ $0.507$ $0.685$ 0 +PTaRL $\mathbf{0.871}$ $\mathbf{0.396}$ $\mathbf{0.722}$ $\mathbf{0.955}$ $0.464$ $\mathbf{0.738}$ 55 +T-JEPA $0.866$ $\pm 2e^{-3}$ $0.351$ $\pm 3e^{-3}$ $0.710$ $\pm 3e^{-3}$ $0.938$ $\pm 4e^{-3}$ $\mathbf{0.448}$ $\pm 2e^{-2}$ $0.517$ $\pm 8e^{-2}$ 11 FT-Trans $0.827$ $0.363$ $0.675$ $0.913$ $0.486$ $0.689$ 0 +PTaRL $\mathbf{0.871}$ $\mathbf{0.397}$ $\mathbf{0.738}$ $\mathbf{0.970}$ $0.448$ $\mathbf{0.738}$ 55 +T-JEPA $0.864$ $\pm 1e^{-3}$ $0.384$ $\pm 5e^{-3}$ $0.708$ $\pm 5e^{-3}$ $0.921$ $\pm 1e^{-2}$ $\mathbf{0.444}$ $\pm 0.131$ $0.551$ $\pm 6e^{-2}$ 11 Other self-supervised methods SwitchTab $0.867$ $0.387$ $0.726$ $0.942$ $0.452$ $0.724$ N/A BinRecon $0.846$ $0.365$ $0.663$ $0.949$ $0.619$ $0.682$ N/A VIME $0.859$ $0.363$ $0.680$ $0.938$ $0.467$ $0.693$ N/A Gradient Boosted Decision Trees (GBDT) XGBoost $\underline{\mathbf{0.874}}$ $0.368$ $0.720$ $0.951$ $0.462$ $\underline{\mathbf{0.729}}$ N/A CatBoost $0.873$ $0.381$ $0.721$ $0.946$ $0.430$ $0.726$ N/A
In Table 6, we present performance metrics for several models, including neural networks (MLP, DCNv2, ResNet, AutoInt, FT-Trans), self-supervised learning (SSL) methods, and gradient-boosted decision trees (GBDT). Regarding T-JEPA, we include standard deviations from 20 independent runs to ensure statistical robustness. Performance metrics for PTaRL [Ye et al., 2024], SwitchTab [Wu et al., 2024], and BinRecon [Lee et al., 2024] were primarily extracted from the original papers, ensuring similar experimental setups for consistency. For VIME, we used the official code made available online by the authors to run the experiments for HE, JA, AL, and HI datasets and report the metrics in their paper for AD and CA.
T-JEPA and PTaRL offer distinct approaches to SSL. T-JEPA focuses on learning robust joint embeddings by predicting future token states, thus enhancing latent space with high-level patterns. It consistently improves deep model performance (e.g., MLP, DCNv2, ResNet), especially in accuracy-based metrics like HE and JA. In contrast, PTaRL learns discrete prototypes from tabular data, capturing key distributional patterns that augment downstream models. PTaRL excels in settings with limited data but lacks consistent statistical variance reporting, limiting direct comparison with T-JEPA. Finally, while PTaRL displays interesting performance across datasets, their approach suffers from a significant computational complexity compared to ours.
We also compare T-JEPA to other SSL methods, such as SwitchTab [Wu et al., 2024], BinRecon [Lee et al., 2024], and VIME [Yoon et al., 2020]. Overall we observe that most approaches augmented by T-JEPA consistently outperform competing SSL approaches apart from PTaRL on some datasets. Nevertheless, we observe that T-JEPA appears to struggle on the HI dataset compared to the existing SSL method, which tend to perform better.

### v3

@@ Appendix C Compute and training time
AD HI HE JA AL CA MNIST Samples 48,842 98,050 65,196 83,733 108,000 20,640 67112 Numerical, Categorical $6$, 88 $28$, 0 $27$, 0 $54$, 0 $128$, 0 88, 0 $784$, 0 Classes 22 22 $100$ 44 $1,000$ N/A $10$ Metric Accuracy Accuracy Accuracy Accuracy Accuracy RMSE Accuracy Pretraining GPU-hours $0.84$ $1.80$ $1.03$ $1.27$ $3.64$ $0.34$ $4.80$
Appendix D Alternative Strategies
Other settings have been considered before the one discussed in section 3, and were discarded because they led to collapsed regimes. We considered two alternatives. First, we considered a pipeline where only the masking strategy differs, as detailed in section D.1. Second, we also considered an alternative where the masking strategy is the one detailed in section D.1, and we also modify the predictor architecture as detailed in section D.2.
D.1 Masking Strategy
@@
Given this modification, the obtained context representation’s dimension differ from the one given in equation (2),
$h_{\text{context}}^{\mathbf{m}}=f_{\theta}(\mathbf{z}_{\mathbf{x}}^{\mathbf{m}})\in\mathbb{R}^{d\times h}.\quad(\text{context})$
The rest of the pipeline is identical to the one given in section 3. We also considered a second alternative where the above pipeline is chosen but a different predictor architecture (see section D.2) replaces the transformer as detailed in section 3.
D.2 Predictor
@@
Appendix E Embedding feature variance
@@
Appendix F Experimental Setting
In Table 1, we present performance metrics for several models, including neural networks (MLP, DCNv2, ResNet, AutoInt, FT-Trans), and with and without T-JEPA. For both cases, we relied on bayesian optimization to find the best set of hyperparameters according to performance on a validation set. We report the performance obtained with this set of hyperparameters on a test set never used during training.
In Table 2 we report the performance of several SSL methods (PTaRL [Ye et al., 2024], SwitchTab [Wu et al., 2024], BinRecon [Lee et al., 2024] and SubTab [Ucar et al., 2021]). Metrics displayed in this table are obtained as follows:
  • For PTaRL [Ye et al., 2024], we report the metrics from their paper for datasets AD, JA and CA as the authors kindly shared the standard deviations corresponding to these datasets for this model. For the remaining datasets (HE, AL, HI and MNIST) we display metrics obtained from our own experiments and their corresponding standard deviations.
  • For BinRecon [Lee et al., 2024], metrics for MLP are primarly obtained from their paper, while we run the experiments ourselves for the ResNet alternative.
  • For SwitchTab [Wu et al., 2024], metrics are obtained from their paper for the MLP dosnwtream model, except for MNIST as they did not use this dataset in their benchmark. Moreover, as the authors have not released any code for their method, we are unable to obtain any metric for the MNIST dataset and the ResNet downstream model. Moreover, no standard deviations are reported in their paper.
  • For VIME [Yoon et al., 2020], we used the official code made available online by the authors to run the experiments for HE, JA, AL, CA and HI datasets and report the metrics from their paper for AD and MNIST for the MLP downstream model, and run the experiments for all datasets for the ResNet downstream model. Notably, we rely on the self-supervised only set-up for fair comparison with other methods.
  • For SubTab [Ucar et al., 2021], we rely on their official implementation and run experiments on each of the datasets except for AD which we obtained from their paper for the MLP downstream model.
For each SSL model, downstream model and dataset we rely on bayesian optimization to set the hyperparameters. This ensure fair and reproducible comparison between all methods. We provide the code to replicate all our experiments on the official github repository and we display in appendix F.1 and F.2 the experimental details.
F.1 Baseline performance
The default hyperparameters for each model were utilized to ensure consistent configurations and establish baseline performance metrics. To maintain alignment with the experimental framework presented in Ye et al. [2024], the same hyperparameters were adopted for generating the baseline results.
Table 12: Default Hyperparameters for MLP
Hyperparameter Value
Number of Layers 4
Hidden Dimension 256
Dropout 0.1
Batch Size 128
Learning Rate $1\times 10^{-4}$
Early Stopping Patience 16
Maximum Epochs 200
Categorical Embedding Dim 128
Table 13: Default Hyperparameters for DCNV2
Hyperparameter Value
Hidden Dimension 128
Number of Cross Layers 3
Number of Hidden Layers 7
Cross Dropout 0.1
Hidden Dropout 0.1
Batch Size 128
Learning Rate $1\times 10^{-4}$
Early Stopping Patience 16
Maximum Epochs 200
Categorical Embedding Dim 128
Table 14: Default Hyperparameters for ResNet
Hyperparameter Value
Number of Layers 4
Hidden Dimension 256
Hidden Dropout 0.1
Batch Size 128
Learning Rate $1\times 10^{-4}$
Early Stopping Patience 16
Maximum Epochs 200
Categorical Embedding Dim 128
Table 15: Default Hyperparameters for AutoInt
Hyperparameter Value
Hidden Dimension 192
Number of Layers 3
Number of Heads 8
Attention Dropout 0.1
Residual Dropout 0.1
Batch Size 128
Learning Rate $1\times 10^{-4}$
Early Stopping Patience 16
Maximum Epochs 200
Table 16: Default Hyperparameters for FT-Transformer
Hyperparameter Value
Hidden Dimension 192
Number of Layers 3
Number of Heads 8
Attention Dropout 0.1
Residual Dropout 0.0
Batch Size 128
Learning Rate $1\times 10^{-4}$
Early Stopping Patience 16
Maximum Epochs 200
F.2 Self-supervised methods
SwitchTab
SwitchTab is a framework for tabular data representation learning that utilizes anencoder-decoder architecture to decouple features into mutual (shared across samples) and salient (unique to individual samples) representations. The process begins with feature corruption applied to input data to enhance robustness. The corrupted data is then encoded by an encoder, producing feature vectors that are decoupled into mutual and salient components using two projectors. These components are recombined and reconstructed by a decoder, with both recovered and switched outputs contributing to the computation of a reconstruction loss.
The results presented are drawn from the original work [Wu et al., 2024]. Since the code was not available, results for the MNIST dataset were not included.
BinRecon
In the approach presented in [Lee et al., 2024], continuous numerical features are discretized into a fixed number of bins, where each bin represents a range of values defined by quantiles of the training dataset. The task is to predict the bin indices instead of reconstructing the raw values, effectively transforming the problem into a regression or classification task depending on whether the bins are treated as ordinal or categorical. This encourages the encoder to learn representations that capture irregularities and nonlinear dependencies characteristic of tabular data. The loss for BinRecon, when treating bins as ordinal values, is defined as:
$L_{BinRecon}=\frac{1}{N}\sum_{i=1}^{N}\|t_{i}-f_{BinRecon}(z_{i})\|^{2},$
where $t_{i}$ denotes the bin indices of the input features, $z_{i}$ represents the encoder outputs, and $f_{BinRecon}$ is the decoder network.
VIME
The VIME framework is a self-supervised learning method for tabular data that leverages two pretext tasks: feature estimation, which involves reconstructing the original features, and mask estimation, which focuses on predicting the applied binary mask. A masked sample $\tilde{\mathbf{x}}$ is generated as:
$\tilde{\mathbf{x}}=\mathbf{m}\odot\bar{\mathbf{x}}+(1-\mathbf{m})\odot\mathbf{x},$
where $\mathbf{m}$ is a binary mask, and $\bar{\mathbf{x}}$ are values sampled from marginal distributions. An encoder $e$ maps $\tilde{\mathbf{x}}$ to $\mathbf{z}$, which is used to predict the mask ($\hat{\mathbf{m}}$) and reconstruct the input ($\hat{\mathbf{x}}$). The framework minimizes:
[…]
where $l_{m}$ is binary cross-entropy and $l_{r}$ is reconstruction loss.
The datasets AD and MNIST were sourced from the work presented in [Yoon et al., 2020]. For the remaining datasets, the publicly available code repositories were utilized, ensuring reproducibility and adherence to standardized evaluation protocols.
The hyperparameters presented in Table 17 were obtained following a hyperparameter tuning process, designed to optimize performance metrics for each specific dataset. This tuning involved Bayesian hyper-parameter search across a range of values for parameters such as alpha, beta, mlp_hidden_dim, and p_m, ensuring that the reported results reflect the best possible configurations for the proposed approach.
Table 17: Hyperparameter Settings for VIME experiments
Dataset alpha beta mlp_hidden_dim p_m
JA $3.1343$ $1.5921\times 10^{0}$ 32 $1.5536\times 10^{-1}$
AL $1.3867$ $1.0233\times 10^{0}$ 256 $2.4229\times 10^{-1}$
HE $4.7568$ $8.1190\times 10^{-1}$ 256 $1.2751\times 10^{-1}$
HI $4.7680$ $1.4418\times 10^{0}$ 128 $2.4642\times 10^{-1}$
CA $2.5088$ 256 $1.1916\times 10^{-1}$
SubTab
SubTab is a framework for representation learning on tabular data inspired by cropping in image augmentation. It divides tabular data into subsets of features, processed by a shared encoder-decoder architecture.
The results in Table 1 were derived using the AD and MNIST datasets, sourced from [Ucar et al., 2021]. For the other datasets, publicly available code repositories were employed, ensuring reproducibility.
The hyperparameters listed in Table 18 were determined through a hyperparameter tuning process aimed at optimizing performance metrics for each dataset. This process utilized Bayesian hyperparameter search across a predefined range of parameter values.
Table 18: Hyperparameter Settings for SubTab experiments
Dataset Dropout Rate Hidden Layers Learning Rate Masking Ratio N Subsets
HE $0.1941$ $[1024,256]$ $1.0867\times 10^{-3}$ $0.2$ 44
JA $0.1383$ $[1024,1024,128]$ $1.6595\times 10^{-3}$ $0.3$ 4
AL $0.1542$ $[1024,512]$ $6.1929\times 10^{-4}$ $0.1$ 66
CA $0.0292$ $[128,128]$ $5.8552\times 10^{-4}$ $0.1$ 44
HI $0.1310$ $[512,512]$ $1.4141\times 10^{-3}$ $0.2$ 44
Appendix G Representation Space Characterization
Kullback-Leibler divergence
$D_{\text{KL}}(P\parallel Q)=\sum_{x}P(x)\log\frac{P(x)}{Q(x)}.$ (8)
Uniformity
We rely on the uniformity score [Wang and Isola, 2020] to evaluate the preservation of maximal information within the feature distribution. This score leverages the Gaussian potential kernel $G_{t}:\mathcal{S}^{d}\times\mathcal{S}^{d}\rightarrow\mathbb{R}_{+}$, defined as $G_{t}(u,v)\triangleq e^{-t\|u-v\|^{2}_{2}},\;t>0$. The uniformity score is defined as the logarithm of the average pairwise Gaussian potential, formally:
Appendix H Regularization Token
[…]
Figure 7: Regularization Token Pipeline. In step (a) a sample $\mathbf{x}\in\mathbb{R}^{d}$ is pre-processed and masked according to a context mask, a [REG] token is appended to the masked representation. In step (b) the whole unmasked sample is pre-processed and fed to the target encoder, including the [REG] token. In step (c) the output of the target encoder is masked according to a target mask (see equation 4). In step (d), the [REG] token is discarded and the remaining features’ representations are fed to the predictor to predict the target output. In step (e) we compute the $\ell_{2}$-distance between the masked target representation and the corresponding prediction.
As discussed in sections 3 and 5, including regularization tokens in the T-JEPA training pipeline is instrumental to escaping representation collapse. While EMA between the weights of the context and target encoders is sufficient for JEPA-based approaches for images [Assran et al., 2023] or videos [Bardes et al., 2024], we show in section 5.2 that without [REG] tokens, the training loss is stuck in a collapsed equilibrium.
We provide in this section a more detailed description on how [REG] are involved in the training pipeline. In particular, we provide in Figure 7 a descriptive example on how the regularization tokens are handled.
[REG] Token Pipeline
Recall that $\mathbf{z}_{\mathbf{x}}^{\mathbf{m}}\in\mathbb{R}^{l_{\mathbf{m}}\times h}$ denotes the obtained embedded representation of sample $\mathbf{x}$ with mask $\mathbf{m}$.
  • Context: Let $\mathbf{m}_{c}$ denote the context mask. The context representation is obtained by concatenating the masked embedded representation and the embedded [REG] token, $\tilde{\mathbf{z}}_{\mathbf{x}}^{\mathbf{m}_{c}}=\texttt{concat}(\mathbf{z}_{\mathbf{x}}^{\mathbf{m}_{c}},\texttt{[REG]})$, and passing it through the context encoder. One obtains
    $\tilde{h}_{\text{context}}^{\mathbf{m}_{c}}=f_{\theta}(\tilde{\mathbf{z}}_{\mathbf{x}}^{\mathbf{m}_{c}})\in\mathbb{R}^{(l_{\mathbf{m}_{c}}+1)\times h},$
    where $l_{\mathbf{m}_{c}}$ corresponds to the number of unmasked features in mask $\mathbf{m}_{c}$, and the supplementary dimension comes from the [REG] token.
  • Target: The target representation is obtained by concatenating the unmasked embedded representation and the embedded [REG] token, $\tilde{\mathbf{z}}_{\mathbf{x}}^{\mathbf{0}_{d}}=\texttt{concat}(\mathbf{z}_{\mathbf{x}}^{\mathbf{0}_{d}},\texttt{[REG]})$ and passing it through the target encoder. One obtains
    $\tilde{h}_{\text{target}}=f_{\theta}(\tilde{\mathbf{z}}_{\mathbf{x}}^{\mathbf{0}_{d}})\in\mathbb{R}^{(d+1)\times h},$
    Then this representation is masked as detailed in equation 4 using the target mask $\mathbf{m}_{t}$, to obtain $h_{\text{target}}^{\mathbf{m}_{t}}$. Note that the [REG] token is always masked/dropped here. See step (c) in Figure 7.
  • The [REG] token is dropped from the context representation that serves to predict the target representation. One thus obtains $h_{\text{context}}^{\mathbf{m}_{c}}$,
    $\tilde{h}_{\text{context}}^{\mathbf{m}_{c}}\xrightarrow{\text{drop {[REG]}}}h_{\text{context}}^{\mathbf{m}_{c}}\in\mathbb{R}^{l_{\mathbf{m}_{c}}\times h}.$
    See step (d) in Figure 7.
  • Prediction: One then uses $h_{\text{context}}^{\mathbf{m}_{c}}$ to predict $h_{\text{target}}^{\mathbf{m}_{t}}$,
    $\hat{h}_{\text{target}}^{\mathbf{m}_{t}}=g_{\phi}(h_{\text{context}}^{\mathbf{m}_{c}},\mathbf{m}_{t}).$
Note that, while we experiment using more than one [REG] token in section 5.2, the rest of the experiments presented in this work include one [REG] token.